### v1

@@ Abstract
Class imbalance remains a major challenge in machine learning, especially in multi-class problems with long-tailed distributions. Existing methods, such as data resampling, cost-sensitive techniques, and logistic loss modifications, though popular and often effective, lack solid theoretical foundations. As an example, we demonstrate that cost-sensitive methods are not Bayes consistent. This paper introduces a novel theoretical framework for analyzing generalization in imbalanced classification. We propose a new class-imbalanced margin loss function for both binary and multi-class settings, prove its strong ${\mathscr{H}}$-consistency, and derive corresponding learning guarantees based on empirical loss and a new notion of class-sensitive Rademacher complexity. Leveraging these theoretical results, we devise novel and general learning algorithms, immax (Imbalanced Margin Maximization), which incorporate confidence margins and are applicable to various hypothesis sets. While our focus is theoretical, we also present extensive empirical results demonstrating the effectiveness of our algorithms compared to existing baselines.
@@ 1 Introduction
The class imbalance problem, defined by a significant disparity in the number of instances across classes within a dataset, is a common challenge in machine learning applications (Lewis and Gale, 1994; Fawcett and Provost, 1996; Kubat and Matwin, 1997; Kang et al., 2021; Menon et al., 2021; Liu et al., 2019; Cui et al., 2019). This issue is prevalent in many real-world binary classification scenarios, and arguably even more so in multi-class problems with numerous classes. In such cases, a few majority classes often dominate the dataset, leading to a “long-tailed” distribution. Classifiers trained on these imbalanced datasets often struggle on the minority classes, performing similarly to a naive baseline that simply predicts the majority class.
The problem has been widely studied in the literature (Cardie and Nowe, 1997; Kubat and Matwin, 1997; Chawla et al., 2002; He and Garcia, 2009; Wallace et al., 2011). While a comprehensive review is beyond our scope, we summarize key strategies into broad categories and refer readers to a recent survey by Zhang et al. (2023) for further details. The primary approaches include the following.
Data modification methods. Techniques such as oversampling the minority classes (Chawla et al., 2002), undersampling the majority classes (Wallace et al., 2011; Kubat and Matwin, 1997), or generating synthetic samples (e.g., SMOTE (Chawla et al., 2002; Qiao and Liu, 2008; Han et al., 2005)), aim to rebalance the dataset before training (Chawla et al., 2002; Estabrooks et al., 2004; Liu et al., 2008; Zhang and Pfister, 2021).
Cost-sensitive techniques. These assign different penalization costs to losses for different classes. They include cost-sensitive SVM (Iranmehr et al., 2019; Masnadi-Shirazi and Vasconcelos, 2010) and other cost-sensitive methods (Elkan, 2001; Zhou and Liu, 2005; Zhao et al., 2018; Zhang et al., 2018, 2019; Sun et al., 2007; Fan et al., 2017; Jamal et al., 2020). The weights are often determined by the relative number of samples in each class or a notion of effective sample size Cui et al. (2019).
These two approaches are closely related and can be equivalent in the limit, with cost-sensitive methods offering a more efficient and principled implementation of data sampling. However, both approaches act by effectively modifying the underlying distribution and risk overfitting minority classes, discarding majority class information, and inherently biasing the training distribution. Very importantly, these techniques may lead to Bayes inconsistency (proven in Section 6). So while effective in some cases, their performance depends on the problem, data distribution, predictors, and evaluation metrics (Van Hulse et al., 2007), and they often require extensive hyperparameter tuning. Hybrid approaches aim to combine these two techniques but inherit many of their limitations.
@@
Other methods. Additional approaches for addressing imbalanced data (see (Zhang et al., 2023)) include post-hoc adjustments of decision thresholds (Fawcett and Provost, 1996; Collell et al., 2016) or class weights (Kang et al., 2020; Kim and Kim, 2019), and techniques like transfer learning, data augmentation, and distillation (Li et al., 2024b).
Despite the many significant advances, these techniques continue to face persistent challenges. Most existing solutions are heuristic-driven and lack a solid theoretical foundation, making their performance unpredictable across diverse contexts. To our knowledge, only Cao et al. (2019) provides an analysis of generalization guarantees, which is limited to the balanced loss, the uniform average of misclassification errors across classes. Their analysis also applies only to binary classification under the separable case and does not address the target misclassification loss.
@@ 2 Preliminaries
Binary classification. Let ${\mathscr{X}}$ represent the input space, and ${\mathscr{Y}}=\left\{-1,+1\right\}$ the binary label space. Let ${\mathscr{D}}$ be a distribution over ${\mathscr{X}}\times{\mathscr{Y}}$, and ${\mathscr{H}}$ a hypothesis set of functions mapping from ${\mathscr{X}}$ to $\mathbb{R}$. Denote by ${\mathscr{H}}_{\mathrm{all}}$ the set of all measurable functions, and by $\ell\colon{\mathscr{H}}_{\mathrm{all}}\times{\mathscr{X}}\times{\mathscr{Y}}\to\mathbb{R}$ a loss function. The generalization error of a hypothesis $h\in{\mathscr{H}}$ and the best-in-class generalization error of ${\mathscr{H}}$ for a loss function $\ell$ are defined as follows: ${\mathscr{R}}_{\ell}(h)=\operatorname*{\mathbb{E}}_{(x,y)\sim{\mathscr{D}}}\left[\ell(h,x,y)\right]$, and ${\mathscr{R}}_{\ell}^{*}({\mathscr{H}})=\inf_{h\in{\mathscr{H}}}{\mathscr{R}}_{\ell}(h)$. The target loss function in binary classification is the zero-one loss function defined for all $h\in{\mathscr{H}}$ and $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$ by $\ell_{0-1}(h,x,y)\coloneqq\mathds{1}_{\operatorname{sign}(h(x))\neq y}$, where $\operatorname{sign}(\alpha)=\mathds{1}_{\alpha\geq 0}-\mathds{1}_{\alpha<0}$. For a labeled example $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$, the margin $\rho_{h}(x,y)$ of a predictor $h\in{\mathscr{H}}$ is defined by $\rho_{h}(x,y)=yh(x)$.
Consistency. A fundamental property of a surrogate loss $\ell_{A}$ for a target loss function $\ell_{B}$ is its Bayes-consistency. Specifically, if a sequence of predictors $\{h_{n}\}_{n\in\mathbb{N}}\subset{\mathscr{H}}_{\rm{all}}$ achieves the optimal $\ell_{A}$-loss asymptotically, then it also achieves the optimal $\ell_{B}$-loss in the limit: $\lim_{n\to+\infty}{\mathscr{R}}_{\ell_{A}}(h_{n})={\mathscr{R}}^{*}_{\ell_{A}}({\mathscr{H}}_{\rm{all}})\Rightarrow\lim_{n\to+\infty}{\mathscr{R}}_{\ell_{B}}(h_{n})={\mathscr{R}}^{*}_{\ell_{B}}({\mathscr{H}}_{\rm{all}})$. While Bayes-consistency is a natural and desirable property, it is inherently asymptotic and applies only to the family of all measurable functions ${\mathscr{H}}_{\rm{all}}$. A more applicable and informative notion is that of ${\mathscr{H}}$-consistent bounds, which account for the specific hypothesis class ${\mathscr{H}}$ and provide non-asymptotic guarantees (Awasthi et al., 2022a, b; Mao et al., 2023f) (see also (Awasthi et al., 2021a, b, 2023, 2024; Mao et al., 2023b, c, d, e, a, 2024c, 2024b, 2024a, 2024e, 2024h, 2024i, 2024d, 2024f, 2024g; Mohri et al., 2024; Cortes et al., 2024)). In the realizable setting, these bounds are of the form:
@@
where $\Gamma$ is a non-increasing concave function with $\Gamma(0)=0$. In the general non-realizable setting, each side of the bound is augmented with a minimizabily gap
${\mathscr{M}}_{\ell}({\mathscr{H}})={\mathscr{R}}_{\ell}^{*}({\mathscr{H}})-\operatorname*{\mathbb{E}}_{x}\left[\inf_{h\in{\mathscr{H}}}\operatorname*{\mathbb{E}}_{y}\left[\ell(h,x,y)\mid x\right]\right],$
which measures the difference between the best-in-class error and the expected best-in-class conditional error. The resulting bound is:
${\mathscr{R}}_{\ell_{B}}(h)-{\mathscr{R}}^{*}_{\ell_{B}}({\mathscr{H}})+{\mathscr{M}}_{\ell_{B}}({\mathscr{H}})\leq\Gamma\left({\mathscr{R}}_{\ell_{A}}(h)-{\mathscr{R}}^{*}_{\ell_{A}}({\mathscr{H}})+{\mathscr{M}}_{\ell_{A}}({\mathscr{H}})\right).$
${\mathscr{H}}$-consistency bounds imply Bayes-consistency when ${\mathscr{H}}={\mathscr{H}}_{\rm{all}}$ (Mao et al., 2024i) and provide stronger and more applicable guarantees.
@@ 3 Theoretical Analysis of Imbalanced Binary Classification
Our theoretical analysis addresses imbalance by introducing distinct confidence margins for positive and negative points. This allows us to explicitly account for the effects of class imbalance. We begin by defining a general class-imbalanced margin loss function based on these confidence margins. Subsequently, we prove that, unlike previously studied cost-sensitive loss functions in the literature, this new loss function satisfies ${\mathscr{H}}$-consistency bounds. Furthermore, we establish general margin bounds for imbalanced binary classification in terms of the proposed class-imbalanced margin loss. While our use of margins bears some resemblance to the interesting approach of Cao et al. (2019), their analysis is limited to geometric margins in the separable case, making ours fundamentally distinct.
@@ 3.2 ${\mathscr{H}}$-Consistency
The proof is presented in Appendix D.2. The next section presents generalization bounds based on the empirical class-imbalanced margin loss, along with the $(\rho_{+},\rho_{-})$-class-sensitive Rademacher complexity and its empirical counterpart defined below. Given a sample $S=\left(x_{1},\ldots,x_{m}\right)$, we define $I_{+}=\left\{i\in\left\{1,\ldots,m\right\}\mid y_{i}=+1\right\}$ and $m_{+}=|I_{+}|$ as the number of positive instances. Similarly, we define $I_{-}=\left\{i\in\left\{1,\ldots,m\right\}\mid y_{i}=-1\right\}$ and $m_{-}=|I_{-}|$ as the number of negative instances.
@@ 4 Algorithm for Binary Classification
$\min_{h\in\overline{\mathscr{H}}}\lambda\left\|h\right\|^{2}+\frac{1}{m}\left[\sum_{i\in I_{+}}\Psi\left(\frac{h(x_{i})}{\alpha}\right)+\sum_{i\in I_{-}}\Psi\left(\frac{-h(x_{i})}{1-\alpha}\right)\right],$
@@
Finally, note that, while $\alpha$ can be freely searched over a range of values in our general (non-separable case) algorithm, it can be beneficial to focus the search around the optimal values identified in the separable case.
@@ 5 Extension to Multi-Class Classification
We will adopt the same notation and definitions as previously described, with some slight adjustments. In particular, we denote the multi-class label space by ${\mathscr{Y}}=[c]\coloneqq\left\{1,\ldots,c\right\}$ and a hypothesis set of functions mapping from ${\mathscr{X}}\times{\mathscr{Y}}$ to $\mathbb{R}$ by ${\mathscr{H}}$. For a hypothesis $h\in{\mathscr{H}}$, the label ${\sf h}(x)$ assigned to $x\in{\mathscr{X}}$ is the one with the largest score, defined as ${\sf h}(x)=\operatorname*{argmax}_{y\in{\mathscr{Y}}}h(x,y)$, using the highest index for tie-breaking. For a labeled example $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$, the margin $\rho_{h}(x,y)$ of a hypothesis $h\in{\mathscr{H}}$ is given by $\rho_{h}(x,y)=h(x,y)-\max_{y^{\prime}\neq y}h(x,y^{\prime})$, which is the difference between the score assigned to $(x,y)$ and that of the next-highest scoring label. We define the multi-class zero-one loss function as $\ell^{\rm{multi}}_{0-1}\coloneqq\mathds{1}_{{\sf h}(x)\neq y}$. This is the target loss of interest in multi-class classification.
@@
The main margin bounds in this section are expressed in terms of this loss function. The parameters $\rho_{k}>0$, for $k\in[c]$, represent the confidence margins imposed by a hypothesis $h$ for instances labeled $k$. As in the binary case, we establish an equivalent expression for this class-imbalanced margin loss function (Lemma E.2). We also prove that our multi-class class-imbalanced ${\boldsymbol{\rho}}$-margin loss is ${\mathscr{H}}$-consistent for any complete hypothesis set ${\mathscr{H}}$ (Theorem E.3). This covers all commonly used function classes in practice, such as linear classifiers and neural network architectures.
Our generalization bounds are expressed in terms of the following notions of ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity.
@@
where $\lambda$ and $\rho_{k}$s are chosen via cross-validation. Let $\rho=\sum_{k=1}^{c}\rho_{k}$ and $\overline{r}=\left[\sum_{k=1}^{c}m_{k}^{\frac{1}{3}}r_{k,2}^{\frac{2}{3}}\right]^{\frac{3}{2}}$. Using Lemma F.4 (Appendix F.4), the term under the square root in the second term of the generalization bound can be reformulated in terms of the Rényi divergence of order 3 as: $\sum_{k=1}^{c}\frac{m_{k}r_{k,2}^{2}}{\rho_{k}^{2}}=\frac{\overline{r}^{2}}{\rho^{2}}e^{2{\mathsf{D}}_{3}\left({\mathsf{r}}\,\|\,\frac{{\boldsymbol{\rho}}}{\rho}\right)}$, where ${\mathsf{r}}=\bigg{[}\frac{m_{k}^{\frac{1}{3}}r_{k,2}^{\frac{2}{3}}}{\overline{r}^{\frac{2}{3}}}\bigg{]}_{k}$. Thus, while $\rho_{k}$s can be freely searched over a range of values in our general algorithm, it may be beneficial to focus the search for the vector $[\rho_{k}/\rho]_{k}$ near ${\mathsf{r}}$. When the number of classes $c$ is very large, the search space can also be significantly reduced by assigning identical $\rho_{k}$ values to underrepresented classes while reserving distinct $\rho_{k}$ values for the most frequently occurring classes.

### v2

@@ Abstract
Class imbalance remains a major challenge in machine learning, especially in multi-class problems with long-tailed distributions. Existing methods, such as data resampling, cost-sensitive techniques, and logistic loss modifications, though popular and often effective, lack solid theoretical foundations. As an example, we demonstrate that cost-sensitive methods are not Bayes-consistent. This paper introduces a novel theoretical framework for analyzing generalization in imbalanced classification. We propose a new class-imbalanced margin loss function for both binary and multi-class settings, prove its strong ${\mathscr{H}}$-consistency, and derive corresponding learning guarantees based on empirical loss and a new notion of class-sensitive Rademacher complexity. Leveraging these theoretical results, we devise novel and general learning algorithms, immax (Imbalanced Margin Maximization), which incorporate confidence margins and are applicable to various hypothesis sets. While our focus is theoretical, we also present extensive empirical results demonstrating the effectiveness of our algorithms compared to existing baselines.
imbalanced data, consistency, margin bounds, learning theory
@@ 1 Introduction
The class imbalance problem, defined by a significant disparity in the number of instances across classes within a dataset, is a common challenge in machine learning applications (Lewis & Gale, 1994; Fawcett & Provost, 1996; Kubat & Matwin, 1997; Kang et al., 2021; Menon et al., 2021; Liu et al., 2019; Cui et al., 2019). This issue is prevalent in many real-world binary classification scenarios, and arguably even more so in multi-class problems with numerous classes. In such cases, a few majority classes often dominate the dataset, leading to a “long-tailed” distribution. Classifiers trained on these imbalanced datasets often struggle on the minority classes, performing similarly to a naive baseline that simply predicts the majority class.
The problem has been widely studied in the literature (Cardie & Nowe, 1997; Kubat & Matwin, 1997; Chawla et al., 2002; He & Garcia, 2009; Wallace et al., 2011). While a comprehensive review is beyond our scope, we summarize key strategies into broad categories and refer readers to a recent survey by Zhang et al. (2023) for further details. The primary approaches include the following.
Data modification methods. Techniques such as oversampling the minority classes (Chawla et al., 2002), undersampling the majority classes (Wallace et al., 2011; Kubat & Matwin, 1997), or generating synthetic samples (e.g., SMOTE (Chawla et al., 2002; Qiao & Liu, 2008; Han et al., 2005)), aim to rebalance the dataset before training (Chawla et al., 2002; Estabrooks et al., 2004; Liu et al., 2008; Zhang & Pfister, 2021).
Cost-sensitive techniques. These assign different penalization costs to losses for different classes. They include cost-sensitive SVM (Iranmehr et al., 2019; Masnadi-Shirazi & Vasconcelos, 2010) and other cost-sensitive methods (Elkan, 2001; Zhou & Liu, 2005; Zhao et al., 2018; Zhang et al., 2018, 2019; Sun et al., 2007; Fan et al., 2017; Jamal et al., 2020). The weights are often determined by the relative number of samples in each class or a notion of effective sample size (Cui et al., 2019).
These two approaches are closely related and can be equivalent in the limit, with cost-sensitive methods offering a more efficient and principled implementation of data sampling. However, both approaches act by effectively modifying the underlying distribution and risk overfitting minority classes, discarding majority class information, and inherently biasing the training distribution. Very importantly, these techniques may lead to Bayes-inconsistency (proven in Section 6). So while effective in some cases, their performance depends on the problem, data distribution, predictors, and evaluation metrics (Van Hulse et al., 2007), and they often require extensive hyperparameter tuning. Hybrid approaches aim to combine these two techniques but inherit many of their limitations.
@@
Other methods. Additional approaches for addressing imbalanced data (see (Zhang et al., 2023)) include post-hoc adjustments of decision thresholds (Fawcett & Provost, 1996; Collell et al., 2016) or class weights (Kang et al., 2020; Kim & Kim, 2020), and techniques like transfer learning, data augmentation, and distillation (Li et al., 2024b).
Despite the many significant advances, these techniques continue to face persistent challenges. Most existing solutions are heuristic-driven and lack a solid theoretical foundation, making their performance unpredictable across diverse contexts. Among prior work, the most closely related is that of (Cao et al., 2019), which provides an analysis of generalization guarantees for the balanced loss, which equalizes the impact of each class by weighting errors inversely to class frequency. Their analysis also applies only to binary classification under the separable case and does not address the target misclassification loss. Jiawei et al. (2020) adopt classical margin theory to derive generalization bounds for multi-class softmax regression, while Wang et al. (2023) establish fine-grained generalization bounds based on data-dependent contraction. Despite these advances, all of these works focus exclusively on the balanced loss. In contrast, our work establishes generalization guarantees with respect to the standard zero-one misclassification loss.
@@ 2 Preliminaries
Binary classification. Let ${\mathscr{X}}$ represent the input space, and ${\mathscr{Y}}=\left\{-1,+1\right\}$ the binary label space. Let ${\mathscr{D}}$ be a distribution over ${\mathscr{X}}\times{\mathscr{Y}}$, and ${\mathscr{H}}$ a hypothesis set of functions mapping from ${\mathscr{X}}$ to $\mathbb{R}$. Denote by ${\mathscr{H}}_{\mathrm{all}}$ the set of all measurable functions, and by $\ell\colon{\mathscr{H}}_{\mathrm{all}}\times{\mathscr{X}}\times{\mathscr{Y}}\to\mathbb{R}$ a loss function. The generalization error of a hypothesis $h\in{\mathscr{H}}$ and the best-in-class generalization error of ${\mathscr{H}}$ for a loss function $\ell$ are defined as follows: ${\mathscr{R}}_{\ell}(h)=\operatorname*{\mathbb{E}}_{(x,y)\sim{\mathscr{D}}}\left[\ell(h,x,y)\right]$, and ${\mathscr{R}}_{\ell}^{*}({\mathscr{H}})=\inf_{h\in{\mathscr{H}}}{\mathscr{R}}_{\ell}(h)$. The target loss function in binary classification is the zero-one loss function defined for all $h\in{\mathscr{H}}$ and $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$ by $\ell_{0-1}(h,x,y)\coloneqq 1_{\operatorname{sign}(h(x))\neq y}$, where $\operatorname{sign}(\alpha)=1_{\alpha\geq 0}-1_{\alpha<0}$. For a labeled example $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$, the margin $\rho_{h}(x,y)$ of a predictor $h\in{\mathscr{H}}$ is defined by $\rho_{h}(x,y)=yh(x)$.
Consistency. A fundamental property of a surrogate loss $\ell_{A}$ for a target loss function $\ell_{B}$ is its Bayes-consistency (Zhang, 2004; Bartlett et al., 2006; Chen et al., 2004). Specifically, if a sequence of predictors $\{h_{n}\}_{n\in\mathbb{N}}\subset{\mathscr{H}}_{\rm{all}}$ achieves the optimal $\ell_{A}$-loss asymptotically, then it also achieves the optimal $\ell_{B}$-loss in the limit: $\lim_{n\to+\infty}{\mathscr{R}}_{\ell_{A}}(h_{n})={\mathscr{R}}^{*}_{\ell_{A}}({\mathscr{H}}_{\rm{all}})\Rightarrow\lim_{n\to+\infty}{\mathscr{R}}_{\ell_{B}}(h_{n})={\mathscr{R}}^{*}_{\ell_{B}}({\mathscr{H}}_{\rm{all}})$. While Bayes-consistency is a natural and desirable property, it is inherently asymptotic and applies only to the family of all measurable functions ${\mathscr{H}}_{\rm{all}}$. A more applicable and informative notion is that of ${\mathscr{H}}$-consistent bounds, which account for the specific hypothesis class ${\mathscr{H}}$ and provide non-asymptotic guarantees (Awasthi et al., 2022a, b, 2021a, 2021b, 2023a, 2023b; Mao et al., 2023f, c, d, e, a, b; Zheng et al., 2023; Mao et al., 2024c, b, a, e, h, f, g; Mohri et al., 2024; Cortes et al., 2024; Mao et al., 2025a, b; Mao, 2025; Zhong, 2025)). In the realizable setting, these bounds are of the form:
@@
where $\Gamma$ is a non-increasing concave function with $\Gamma(0)=0$. In the general non-realizable setting, each side of the bound is augmented with a minimizabily gap ${\mathscr{M}}_{\ell}({\mathscr{H}})={\mathscr{R}}_{\ell}^{*}({\mathscr{H}})-\operatorname*{\mathbb{E}}_{x}\left[\inf_{h\in{\mathscr{H}}}\operatorname*{\mathbb{E}}_{y}\left[\ell(h,x,y)\mid x\right]\right]$, which measures the difference between the best-in-class error and the expected best-in-class conditional error. The resulting bound is: ${\mathscr{R}}_{\ell_{B}}(h)-{\mathscr{R}}^{*}_{\ell_{B}}({\mathscr{H}})+{\mathscr{M}}_{\ell_{B}}({\mathscr{H}})\leq\Gamma\left({\mathscr{R}}_{\ell_{A}}(h)-{\mathscr{R}}^{*}_{\ell_{A}}({\mathscr{H}})+{\mathscr{M}}_{\ell_{A}}({\mathscr{H}})\right)$. ${\mathscr{H}}$-consistency bounds imply Bayes-consistency when ${\mathscr{H}}={\mathscr{H}}_{\rm{all}}$ (Mao et al., 2024i, d) and provide stronger and more applicable guarantees.
@@ 3 Theoretical Analysis of Imbalanced Binary Classification
Our theoretical analysis addresses imbalance by introducing distinct confidence margins for positive and negative points. This allows us to explicitly account for the effects of class imbalance. We begin by defining a general class-imbalanced margin loss function based on these confidence margins. Subsequently, we prove that, unlike previously studied cost-sensitive loss functions in the literature, this new loss function satisfies ${\mathscr{H}}$-consistency bounds. Furthermore, we establish general margin bounds for imbalanced binary classification in terms of the proposed class-imbalanced margin loss. While our use of margins bears some resemblance to the approach of Cao et al. (2019), their analysis is limited to geometric margins in the separable case, making ours fundamentally distinct.
@@ 3.2 ${\mathscr{H}}$-Consistency
The proof is presented in Appendix D.2. Note that our ${\mathscr{H}}$-consistency bounds in Theorem 3.3 can be extended directly to the uniformly bounded hypothesis sets considered in Theorem 4.1 below. In this case, the bounds would depend on the complexity of the hypothesis class, similar to the ${\mathscr{H}}$-consistency bounds presented in (Awasthi et al., 2022a). The next section presents generalization bounds based on the empirical class-imbalanced margin loss, along with the $(\rho_{+},\rho_{-})$-class-sensitive Rademacher complexity and its empirical counterpart defined below. Given a sample $S=\left(x_{1},\ldots,x_{m}\right)$, we define $I_{+}=\left\{i\in\left\{1,\ldots,m\right\}\mid y_{i}=+1\right\}$ and $m_{+}=|I_{+}|$ as the number of positive instances. Similarly, we define $I_{-}=\left\{i\in\left\{1,\ldots,m\right\}\mid y_{i}=-1\right\}$ and $m_{-}=|I_{-}|$ as the number of negative instances.
@@ 4 Algorithm for Binary Classification
$\min_{h\in\overline{\mathscr{H}}}\lambda\left\|h\right\|^{2}+\frac{1}{m}\left[\sum_{i\in I_{+}}\Psi\left(\frac{h(x_{i})}{\alpha}\right)+\sum_{i\in I_{-}}\Psi\left(\frac{-h(x_{i})}{1-\alpha}\right)\right],$ (2)
@@
Finally, note that, while $\alpha=\frac{\rho_{+}}{\rho_{+}+\rho_{-}}>0$ in the optimization problem (2) can be freely searched over a range of values in our general (non-separable case) algorithm, it can be beneficial to focus the search around the optimal values identified in the separable case.
@@ 5 Extension to Multi-Class Classification
We will adopt the same notation and definitions as previously described, with some slight adjustments. In particular, we denote the multi-class label space by ${\mathscr{Y}}=[c]\coloneqq\left\{1,\ldots,c\right\}$ and a hypothesis set of functions mapping from ${\mathscr{X}}\times{\mathscr{Y}}$ to $\mathbb{R}$ by ${\mathscr{H}}$. For a hypothesis $h\in{\mathscr{H}}$, the label ${\mathsf{h}}(x)$ assigned to $x\in{\mathscr{X}}$ is the one with the largest score, defined as ${\mathsf{h}}(x)=\operatorname*{argmax}_{y\in{\mathscr{Y}}}h(x,y)$, using the highest index for tie-breaking. For a labeled example $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$, the margin $\rho_{h}(x,y)$ of a hypothesis $h\in{\mathscr{H}}$ is given by $\rho_{h}(x,y)=h(x,y)-\max_{y^{\prime}\neq y}h(x,y^{\prime})$, which is the difference between the score assigned to $(x,y)$ and that of the next-highest scoring label. We define the multi-class zero-one loss function as $\ell^{\rm{multi}}_{0-1}\coloneqq 1_{{\mathsf{h}}(x)\neq y}$. This is the target loss of interest in multi-class classification.
@@
The main margin bounds in this section are expressed in terms of this loss function. The parameters $\rho_{k}>0$, for $k\in[c]$, represent the confidence margins imposed by a hypothesis $h$ for instances labeled $k$. As in the binary case, we establish an equivalent expression for this class-imbalanced margin loss function.
Lemma 5.2.
${\mathsf{L}}_{{\boldsymbol{\rho}}}(h,x,y)=\sum_{k=1}^{c}\Phi_{\rho_{k}}\left(\rho_{h}(x,y)\right)1_{{\mathsf{h}}(x)=k}.$
The proof is included in Appendix F.1. We also prove that our multi-class class-imbalanced ${\boldsymbol{\rho}}$-margin loss is ${\mathscr{H}}$-consistent for any complete hypothesis set ${\mathscr{H}}$. We say a hypothesis set is complete when the scoring values spanned by ${\mathscr{H}}$ for each instance cover $\mathbb{R}$: for all $(x,y)\in{\mathscr{X}}\times{\mathscr{Y}}$, $\left\{h(x,y)\colon h\in{\mathscr{H}}\right\}=\mathbb{R}$. This covers all commonly used function classes in practice, such as linear classifiers and neural network architectures.
Theorem 5.3 (${\mathscr{H}}$-Consistency bound for multi-class class-imbalanced margin loss).
The proof is included in Appendix F.2. Our generalization bounds are expressed in terms of the following notions of ${\boldsymbol{\rho}}$-class-sensitive Rademacher complexity.
@@
where $\lambda$ and $\rho_{k}$s are chosen via cross-validation. Let $\rho=\sum_{k=1}^{c}\rho_{k}$ and $\overline{r}=\left[\sum_{k=1}^{c}m_{k}^{\frac{1}{3}}r_{k,2}^{\frac{2}{3}}\right]^{\frac{3}{2}}$, where as defined previously, $r_{k}=\sup_{i\in I_{k},y\in{\mathscr{Y}}}\left\|\Phi(x_{i},y)\right\|$, for any $k\in[c]$. Using Lemma F.1 (Appendix F.4), the term under the square root in the second term of the generalization bound can be reformulated in terms of the Rényi divergence of order 3 as: $\sum_{k=1}^{c}\frac{m_{k}r_{k,2}^{2}}{\rho_{k}^{2}}=\frac{\overline{r}^{2}}{\rho^{2}}e^{2{\mathsf{D}}_{3}\left({\mathsf{r}}\,\|\,\frac{{\boldsymbol{\rho}}}{\rho}\right)}$, where ${\mathsf{r}}=\bigg{[}\frac{m_{k}^{\frac{1}{3}}r_{k,2}^{\frac{2}{3}}}{\overline{r}^{\frac{2}{3}}}\bigg{]}_{k}$. Thus, while $\rho_{k}$s can be freely searched over a range of values in our general algorithm (8), for experimental efficiency it may be beneficial to focus the search for the vector $[\rho_{k}/\rho]_{k}$ near ${\mathsf{r}}$. When the number of classes $c$ is very large, the search space can also be significantly reduced by assigning identical $\rho_{k}$ values to underrepresented classes while reserving distinct $\rho_{k}$ values for the most frequently occurring classes.